# A Digital Twin for Telesurgery under Intermittent Communication

### v1

@@ Abstract
Telesurgery is an effective way to deliver service from expert surgeons to areas without immediate access to specialized resources. However, many of these areas, such as rural districts or battlefields, might be subject to different problems in communication, especially latency and intermittent periods of communication outage. This challenge motivates the use of a digital twin for the surgical system, where a simulation would mirror the robot hardware and surgical environment in the real world. The surgeon would then be able to interact with the digital twin during communication outage, followed by a recovery strategy on the real robot upon reestablishing communication. This paper builds the digital twin for the da Vinci surgical robot, with a buffering and replay strategy that reduces the mean task completion time by 23% when compared to the baseline, for a peg transfer task subject to intermittent communication outage.
@@ I Introduction
Telesurgery, the process of conducting surgeries remotely with the assistance of robotic systems, has been gaining increasing attention. Performing surgical procedures in this remote fashion would especially benefit less privileged areas without convenient access to specialized surgical resources [1]. Telesurgery addresses the issue of the disproportionately low number of expert surgeons compared to the overall population size in low- to medium-income countries [2, 3]. It can also reduce the cost of travel for both patients and medical professionals [4, 5], as well as facilitating the training of prospective and less experienced surgeons [6].
@@ II Related works
While these works present important applications of AR in surgeries, digital twins, like the one presented in this paper, extend far beyond the ability of basic overlays. By mirroring the configuration of the real world in a virtual environment, digital twins can utilize privileged information in simulation to compensate for imperfections in the real setup, which can be particularly crucial in a surgical environment. The imperfection we focus on in our work is intermittent communication outages, and the work by Richter et al. [22] bears the most resemblance to ours, since it instead considers communication latency, also assisted by a virtual environment when teleoperating a da Vinci system. The authors developed a system where a velocity-based kinematic extrapolation is used to display the position of the instrument a short time period into the future as a virtual overlay, with positioning accuracy mediated by an Extended Kalman Filter. Our work draws key distinctions from [22] as: (1) we focus on communication outage, rather than communication latency, and (2) we do not use a geometry-based visualization, but instead build a digital twin based on a physics simulation, which would allow interactions with other objects in the environment to be computed from dynamics, even during periods of communication outage.
Teleoperation subject to varying amounts of latency is closely related to the communication outage problem we are considering, and many methods have been developed for solving latency issues. This is especially common in bilateral teleoperation, which includes haptic feedback to the operator, because latency tends to destabilize bilateral control systems. The most relevant approaches, which can also apply to communication outage, are those that rely on a model of the remote environment, such as model-mediated telemanipulation [14] where the operator directly interacts with a model that is updated by feedback from the remote system. Similarly, in teleprogramming [23] and tele-sensor-programming [24], the operator interacts with a simulated remote environment and teleprograms the remote robot through a sequence of elementary motion commands. Our approach differs in that the operator interacts with the remote environment during normal communication, and the system automatically switches to the digital twin (simulation model) during communication outage.
@@ III-C Control strategies
When the communication state is normal, the input of the MTM is processed into a Cartesian goal, which is sent to both the remote robot and the digital twin. This was easily achieved by ensuring that the digital robot followed the Collaborative Robotics Toolkit (CRTK) convention [29], a standardized set of commands and measurements for robotic systems, such as the servo_cp command observed in panel (a) of Fig. 1. Using CRTK as a common control interface simplifies the simultaneous control of the real and virtual robots and allows using any CRTK-compatible teleoperation device, such as a 6 DoF joystick or a small tabletop robot arm, to control the robots. Code for setting up the digital twin in AMBF and performing simultaneous control can be found in [30].
@@
For the replay strategy, we aimed to provide a smoother experience to the user by switching teleoperation to the digital twin robot during communication outages and then have the remote robot catch up with the digital twin upon restoring communication. In this case, when the communication is interrupted, the remote robot still remains stationary, but the digital twin is displayed as an AR overlay to enable the user to continue the task. During the outage, the user’s commands are only sent to the digital twin and to a buffer as seen in panel (b) of Fig. 1. When communication is restored, the buffer is played back by skipping every other entry, making the remote robot follow the buffered user input at twice the speed (see panel (c) in Fig. 1). Note that the buffer continues to be appended with new commands during the recovery phase. For example, if the communication outage lasted for 1 s, the system would need 0.5 s to finish replaying the existing buffer, but during this time, new user input also enters the buffer. As a result, a total of 1 s is required after the communication is restored to fully deplete the buffer and catch up to the user’s input, before returning to normal teleoperation.
@@ III-D Digital twin implementation on the AMBF simulator
After performing the camera and environment calibration, the digital twin framework was implemented using AMBF as a simulation environment and ROS as a communication channel (see Fig. 3). To implement custom functionality in AMBF, we leveraged its plugin system which enables compiled libraries to be attached to any AMBF object for additional runtime functionalities. Specifically, two plugins were developed for the digital twin, one to enable overlaying the digital twin on the endoscopic images, and the other one to create a CRTK communication interface for the digital twin.
@@
The first plugin, ROS_AR_CAMERA_PLUGIN, was attached to the simulator cameras to render virtual objects on top of endoscopic images. At initialization, this plugin subscribes to a ROS topic containing the calibration parameters and then modifies the projection matrix of the OpenGL camera inside AMBF based on these parameters. At runtime, the plugin receives images from the camera and uses them as a texture that is applied to the virtual world’s background. This enables overlaying the digital twin on the endoscopic view of the user. This plugin assumes that the endoscope has been previously calibrated and that parameters and images are streamed using ROS as described in Sec. III-B. The second plugin, CRTK_PLUGIN, was attached to the simulated robot to create the CRTK interface that enables the virtual robot to be commanded with the same commands as the real robot.
@@ IV-A Experiment setup
Participants performed the task under the two developed control strategies: baseline and replay. In both conditions, participants experienced simulated, random periods of communication outage. Periods of normal communication have a mean of 3.2 seconds with a standard deviation of 0.15 seconds, while periods of communication outage have a mean of 0.8 seconds with a standard deviation of 0.1 seconds. These values were selected based on previous studies [15, 31, 32]. Due to a system issue, we were unable to append commands to the buffer during the recovery phase, so only the buffered inputs from the communication outage period were replayed, but this should not have a significant effect on the participants’ task performance. Each user’s time to fully complete the peg transfer task from the left to the right and then back is recorded, which includes both the duration of normal communication and the duration of communication outage. This task completion time serves as the first metric to compare between the two different conditions.
@@ V Discussion
Similar benefits for using the digital twin can also be observed from the NASA TLX results in Fig. 5. Although the $t$-test does not report significance, we do see a uniform decrease in the survey response numbers with the replay condition compared to the baseline, which indicates a lower task load in general. One of the largest drops is with the “frustration” dimension, which can be attributed to the fact that the user’s motion with the MTM is often disrupted during the baseline condition. On the contrary, the replay condition eliminates this effect and replaces it with a smooth user experience, hence decreasing user frustration while operating the system.
@@ VI Conclusions and Future Work
This paper introduces a system for constructing and registering a digital twin, created in the AMBF simulator, for the da Vinci Research Kit, and the application of this digital twin to provide assistance during communication outage between the master console and the remote robot. The system framework can be seamlessly adapted to any input device through a CRTK interface, so one direction for future work would be to explore the use of alternative devices to the traditional surgeon console to control both the physical and the digital da Vinci PSM. Another topic would be to explore different patterns of communication imperfection, including both time delays and communication failures. The proposed buffer/replay strategy, where the buffer is replayed at twice the original speed, represents a first step toward mitigation of communication outage. A more sophisticated approach would be to vary the speed increase so that slow motions are increased more than fast motions.
Moreover, there are various opportunities to design different recovery strategies. In particular, those that involve autonomous PSM motion during communication outage could be of particular interest, when they are combined with strategies to remedy the discrepancy between the autonomous motion and the user input after reestablishing communication. These autonomous motions can be derived analytically given ground truth knowledge about the task environment. Alternatively, privileged information from the digital twin could benefit the distillation of some teacher-student reinforcement learning policy that can predict short open-loop action sequence given the history of motion up until communication loss, similar to the approach in [33]. For the average length of communication outage considered in this paper, rolling out these trajectories open-loop could be promising in having the remote robot predict user input when there is no communication.
Finally, the peg transfer task does not involve significant interaction with the environment, which could be challenging to simulate in the digital twin, especially when the interaction involves soft tissue. We adopted the AMBF dynamic simulator to provide a platform that could, in the future, be extended to simulate more complex interactions. It is also important to note that even with high-fidelity dynamic simulation during periods of communication outage, perception of the remote environment will be required to update the digital twin when communication is restored.
Acknowledgments and disclosures
The authors would like to express their appreciation for Haoying Zhou, Adnan Munawar, and Ying Yuan for their support and feedback. H. Ishida was supported by NIDCD K08 Grant DC019708, by a research agreement with the Hong Kong Multi-Scale Medical Robotics Centre and by Johns Hopkins University internal funds.

### v2

@@ Abstract
Telesurgery is an effective way to deliver service from expert surgeons to areas without immediate access to specialized resources. However, many of these areas, such as rural districts or battlefields, might be subject to different problems in communication, especially latency and intermittent periods of communication outage. This challenge motivates the use of a digital twin for the surgical system, where a simulation would mirror the robot hardware and surgical environment in the real world. The surgeon would then be able to interact with the digital twin during communication outage, followed by a recovery strategy on the real robot upon reestablishing communication. This paper builds the digital twin for the da Vinci surgical robot, with a buffering and replay strategy that reduces the mean task completion time by 23% when compared to the baseline, for a peg transfer task subject to intermittent communication outage. The relevant code can be found here: https://github.com/LCSR-CIIS/dvrk_digital_twin_teleoperation.
@@ I Introduction
Telesurgery, the process of conducting surgeries remotely with the assistance of robotic systems, has been gaining increasing attention. Performing surgical procedures in this remote fashion would especially benefit less privileged areas without convenient access to specialized surgical resources [1, 2]. Telesurgery addresses the issue of the disproportionately low number of expert surgeons compared to the overall population size in low- to medium-income countries [3, 4]. It can also reduce the cost of travel for both patients and medical professionals [5, 6], as well as facilitating the training of prospective and less experienced surgeons [7].
@@ II Related works
While these works present important applications of AR in surgeries, digital twins, like the one presented in this paper, extend far beyond the ability of basic overlays. By mirroring the configuration of the real world in a virtual environment, digital twins can utilize privileged information in simulation to compensate for imperfections in the real setup, which can be particularly crucial in a surgical environment. The imperfection we focus on in our work is intermittent communication outages in telesurgery.
Multiple studies have reported the presence of time delays in long-distance teleoperation for surgical robotics [23, 24], which is closely related to the communication outage problem under consideration. Lum et al. demonstrated that even a latency of 0.25 to 0.5 seconds can double the task completion time in teleoperated surgery [24]. Teleoperation subject to varying amounts of latency presents challenges analogous to those encountered in communication outages, and many methods have been developed for solving latency issues[25, 26], [26], [27]. This is especially common in bilateral teleoperation, which includes haptic feedback to the operator, because latency tends to destabilize bilateral control systems.
The work by Richter et al. [27] bears the most resemblance to ours, since they use a virtual environment to assist in overcoming communication latency when teleoperating a da Vinci system. The authors developed a system where a velocity-based kinematic extrapolation is used to display the position of the instrument a short time period into the future as a virtual overlay, with positioning accuracy mediated by an Extended Kalman Filter. Our work draws key distinctions from [27] as: (1) we focus on communication outage, rather than communication latency, and (2) we do not use a geometry-based visualization, but instead build a digital twin based on a physics simulation, which would allow interactions with other objects in the environment to be computed from dynamics, even during periods of communication outage.
The most relevant approaches that can also apply to communication outage, are those that rely on a model of the remote environment, such as model-mediated telemanipulation [15] where the operator directly interacts with a model that is updated by feedback from the remote system. Similarly, in teleprogramming [28] and tele-sensor-programming [29], the operator interacts with a simulated remote environment and teleprograms the remote robot through a sequence of elementary motion commands. Our approach differs in that the operator interacts with the remote environment during normal communication, and the system automatically switches to the digital twin (simulation model) during communication outage.
@@ III-C Control strategies
When the communication state is normal, the input of the MTM is processed into a Cartesian goal, which is sent to both the remote robot and the digital twin. This was easily achieved by ensuring that the physical and digital robots followed the Collaborative Robotics Toolkit (CRTK) convention [34], a standardized set of commands and measurements for robotic systems, such as the servo_cp command observed in panel (a) of Fig. 1. Using CRTK as a common control interface simplifies the simultaneous control of the real and virtual robots and allows using any CRTK-compatible teleoperation device, such as a 6 DoF joystick or a small tabletop robot arm, to control the robots. Code for setting up the digital twin in AMBF and performing simultaneous control can be found in [35].
@@
For the replay strategy, we aim to provide a smoother experience to the user by switching teleoperation to the digital twin robot during communication outages and then having the remote robot catch up with the digital twin upon restoring communication. In this case, when the communication is interrupted, the remote robot still remains stationary, but the digital twin is displayed as an AR overlay to enable the user to continue the task. During the outage, the user’s commands are only sent to the digital twin and to a buffer as seen in panel (b) of Fig. 1. When communication is restored, the buffer is played back by skipping every other entry, making the remote robot follow the buffered user input at twice the speed (see panel (c) in Fig. 1). Note that the buffer continues to be appended with new commands during the recovery phase. For example, if the communication outage lasted for 1 s, the system would need 0.5 s to finish replaying the existing buffer, but during this time, new user input also enters the buffer. As a result, a total of 1 s is required after the communication is restored to fully deplete the buffer and catch up to the user’s input, before returning to normal teleoperation. Before the buffer is depleted, the AR overlay of the digital twin is shown, so that the operator would be able to continue their actions while observing the remote robot catch up. We design this recovery strategy specifically targeting relatively static scenes and short communication outages, and a different approach may be necessary in a highly dynamic environment and/or with long communication outages.
@@ III-D Digital twin implementation on the AMBF simulator
After performing the camera and environment calibration, the digital twin framework was implemented using AMBF as a simulation environment and ROS as a communication channel (see Fig. 3). To implement custom functionality in AMBF, we leveraged its plugin system which enables compiled libraries to be attached to any AMBF object for additional runtime functionalities. Specifically, two plugins were developed for the digital twin, one to overlay the virtual robot on the endoscopic images, and the other to create the CRTK communication interface.
@@
The first plugin, ROS_AR_CAMERA_PLUGIN, was attached to the simulator cameras to render virtual objects on top of endoscopic images. At initialization, this plugin subscribes to a ROS topic containing the calibration parameters and then modifies the projection matrix of the OpenGL camera inside AMBF based on these parameters. At runtime, the plugin receives images from the camera and uses them as a texture that is applied to the virtual world’s background. This enables overlaying the digital twin on the endoscopic view that is presented to the user. This plugin assumes that the endoscope has been previously calibrated and that parameters and images are streamed using ROS as described in Sec. III-B. The second plugin, CRTK_PLUGIN, was attached to the simulated robot to create the CRTK interface that enables the virtual robot to be commanded with the same commands as the real robot.
@@ IV-A Experiment setup
Participants performed the task under the two developed control strategies: baseline and replay. In both conditions, participants experienced simulated, random periods of communication outage. Periods of normal communication have a mean of 3.2 seconds with a standard deviation of 0.15 seconds, while periods of communication outage have a mean of 0.8 seconds with a standard deviation of 0.1 seconds. These values were selected based on previous studies [16, 36, 37]. Due to a system issue, we were unable to append commands to the buffer or display the AR overlay during the recovery phase, so only the buffered inputs from the communication outage period were replayed, and without an overlay. These changes should only have a small negative impact on the users’ experience, and thus we expect the results for the complete system to be better than the ones presented in IV-B. Each user’s time to fully complete the peg transfer task from the left to the right and then back is recorded, which includes both the duration of normal communication and the duration of communication outage. This task completion time serves as the first metric to compare between the two different conditions.
@@ V Discussion
Similar benefits for using the digital twin can also be observed from the NASA TLX results in Fig. 5. Although the $t$-test does not report significance, we do see a uniform decrease in the survey response numbers with the replay condition compared to the baseline, which indicates a lower task load in general. One of the largest drops is in the “frustration” dimension, which can be attributed to the fact that the user’s motion with the MTM is often disrupted during the baseline condition. On the contrary, the replay condition eliminates this effect and replaces it with a smooth user experience, hence decreasing user frustration while operating the system.
@@ VI Conclusions and Future Work
This paper introduces a digital twin system for the dVRK robot, and demonstrates its utility to provide assistance during communication outages in the context of telerobotic surgery. One natural extension of our work is to explore different communication imperfection patterns, including time delays and communication failures, and their mitigation strategies. The proposed buffer/replay strategy, where the buffer is replayed at twice the original speed, represents a first step toward mitigation of communication outage. However, mitigating more complex communication failures could require a more sophisticated replay mechanism, such as varying the replay speed so that slow motions are increased more than fast motions.
Alternatively, one could consider mitigation strategies that involve autonomous PSM motion during communication outage, and upon reestablishing communication, remedy the discrepancy between the autonomous motion and the user input. For instance, privileged information from the digital twin could benefit the distillation of some teacher-student reinforcement learning policy that can predict a short open-loop action sequence given the history of motion up until communication loss, similar to the approach in [38]. For the lengths of communication outage considered in this paper, rolling out these trajectories open-loop when there is no communication could be a reliable prediction of user input.
Moreover, our approach was only tested in a relatively static and known environment. To tackle more realistic surgical tasks, it is critically important to model the environment using available sensors such as RGBD cameras. This would enable the system to modify the buffered motions based on physical abnormalities not accurately simulated in the digital twin or adjustments due to motions near high-risk areas of operation. Overall, rather than replaying the raw motions, the system should identify the user’s intended actions during periods of communication outage and adapt buffered motions based on the environment model to ensure safe execution.
Acknowledgments
The authors would like to express their appreciation for Haoying Zhou and Adnan Munawar for their support and feedback. H. Ishida was supported by NIDCD K08 Grant DC019708, by a research agreement with the Hong Kong Multi-Scale Medical Robotics Centre and by Johns Hopkins University internal funds.